You are the manager of a firm that competes against four other firms by bidding for government contracts. While you believe your product is better than the competition, the government purchasing agent views the products as identical and purchases from the firm offering the best price. Total government demand is Q = 900 -5P and all five firms produce at a constant marginal cost of $40. For security reasons, the government has imposed restrictions that permit a maximum of five firms to compete in this market; thus entry by new firms is prohibited. A member of Congress is concerned because no restrictions have been placed on the price that the government pays for this product. In response, she has proposed legislation that would award each existing firm 20 percent of a contract for 650 units at a contracted price of $50 per unit. If this legislation is passed, by how much should you expect your profits to change?
Answer #1

Answer :- Page No - 1 demand as Total government 0 = 900-52 and the constant marginal cost is Me = 40 Now government oustoricPage No - 2 Now. ML = 40 Hence MR-MC amplis 180 - 040040 - 180-40 c0:40 140 = 0.40 140 003350) and price will be p= - 180-0.2Page No - (3) & CPX ) - 14000 = (130X350) - 1000 T = $31,500 Nevol congress proposed a degustation that qorrse should be formPage No ! - (4) our this, if the degislation profst will decline by is pussed, then $ 25,000